Find the instantaneous rate of change of the function f(x) = x2 + 3x + 4 when x = 2 using First Principles. Confirm that your answer is correct using the derivative rules.


1
Expert's answer
2021-02-28T15:46:29-0500

"f(x)=x^2+3x+4"

"a=2"


Formula for instantaneous rate of change is  "lim_{h\rightarrow0}\dfrac{f(a+h)-f(a)}{h}"


"lim_{h\rightarrow0}\dfrac{f(2+h)-f(2)}{h}"


"lim_{h\rightarrow0}\dfrac{[(2+h)^2+3(2+h)+4]-14}{h}"


"lim_{h\rightarrow0}\dfrac{h(h+7)}{h}"

"lim_{h\rightarrow0}(h+7)= 7"


Derivative rule:

"\dfrac{d}{dx}(x^2+3x+4)=2x+3"


at x=2

"f(x)= 2(2)+3=7"


Hence, the answer is confirmed